kirsten can type a 50-page paper in 8 hours. Last month kirsten and courtney,together,typed a 50-page paper in 6 hours. How long would it take courtney to type a 50-page paper on her own?
Answer by richwmiller(17219)

6/8+6/x=1
6/x=2/8
48=2x
x=24
another way to look at
in six hours kirsten can do 6/8 of the job
6/8=3/4
So kirsten is going to be doing 3/4 of the job
and Courtney is only going to do 1/4
so we need a fraction with 6 as the numerator that equals 1/4
6/x=1/4
x=24 hours
Courtney is slower than molasses in January.
She probably makes too many mistakes too.